OpenStudy (anonymous):

Rebekah kicks a soccer ball off the ground and in the air with an initial velocity of 25 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

assuming that it goes straight up in the air, then \(s=0\) and \(v=25\) making your equation \[h(t)=-16t^2+25t\] max is the second coordinate of the vertex the first coordinate is \(-\frac{b}{2a}=-\frac{25}{2\times (-16)}=\frac{25}{32}\) plug that in for \(t\) and see what you get